why x-ray telesscope kept beyond the earth;s atmosphere? which telescope optical or x-ray would have higher resolving power for the same aperture?
1
Expert's answer
2020-02-28T10:35:53-0500

Because the Earth's atmosphere, is opaque to X-rays, X-ray telescopes must be mounted on high altitude rockets, balloons or artificial satellites [1]. Resolving power is the ability of an imaging device to separate (i.e., to see as distinct) points of an object that are located at a small angular distance. Rayleigh defended this criteria on sources of equal strength.[2]

Considering diffraction through a circular aperture, this translates into:

"{\displaystyle \theta =1.220{\frac {\lambda }{D}}}"

where θ is the angular resolution (radians), λ is the wavelength of light, and D is the diameter of the lens' aperture. The factor 1.220 is derived from a precised calculation of the diffraction pattern of circular aperture. As we can see for the same diameter of aperture X-ray telescopes would have higher resolving power, because "\lambda_{xray}<<\lambda_{light}" and "\theta_{xray}<<\theta_{light}".

But there is some difficulty to made X-ray telescope with the same diameter of aperture as optical. Due to the short "\lambda_{xray}" (of the order of 1 - 20 Angstrom) which is close to the distance between atoms in metals and other structural materials, the design of mirrors and lenses are very different and very difficult to implement than for optics "\lambda_{light} \approx1000-8000" Angstrom. The fact that only sliding beams X-ray may be reflected from mirrors [3].



This is the reason Modern X-ray telescopes have achieved a resolution of 0.5 arc second which is much worse than optical ones ("\approx" 0.01arc second for "D=8m" ).

[1]https://en.wikipedia.org/wiki/X-ray_telescope

[2]https://en.wikipedia.org/wiki/Angular_resolution

[3]https://arxiv.org/ftp/arxiv/papers/1010/1010.4892.pdf